OpenStudy (anonymous):

how do i solve f(x)=sec x, (pi/3,2)

myininaya (myininaya):

What do you mean solve f(x)=sec(x), (pi/3,2)?

OpenStudy (anonymous):

its calculus.. product and quotient rules and higher order derivatives

myininaya (myininaya):

I still don't have any instructions what to with it f(x)=sec(x), (pi/3,2) Does the instructions say to differentiate f(x)=sec(x) at (pi/3,2) Does the instructions say to find the tangent line to f(x)=sec(x) at (pi/3,2)

myininaya (myininaya):

how to solve f(x)=sec(x), (pi/3,2) gives no instructions to anyone

myininaya (myininaya):

in fact you cannot solve f(x)=sec(x) , (pi/3,2)

myininaya (myininaya):

you can solve 2=sec(x) for x and one value of x satisfying that is x=pi/3

myininaya (myininaya):

Assuming you want to differentiate f(x)=sec(x) at (pi/3,2) You would first to need to know (sec(x))'=sec(x)tan(x)

OpenStudy (anonymous):

i apologize, the directions are to find the tangent line to that problem

myininaya (myininaya):

Ok lines have the form y-y_1=m(x-x_1) where (x_1,y_1) is a point you know on the line And since we are talking about tangent lines at (pi/3,2) we know (x_1,y_1) is (pi/3,2) since it is on the line m is the slope in order to find the slope you must first find f'(x) then evaluate f'(pi/3)

myininaya (myininaya):

so your equation will look like this: \[y-2=f'(\frac{\pi}{3})(x-\frac{\pi}{3})\]

myininaya (myininaya):

you just have to find the f'(pi/3) part

OpenStudy (anonymous):

so i just have to find the slope of the for the equation ? im sorry im just really lost in calc.

myininaya (myininaya):

yes and slopes can be found from the derivative

OpenStudy (anonymous):

all right

myininaya (myininaya):

so you need to know the derivative of sec(x)

myininaya (myininaya):

which i will give you since it should be something you commit to memory anyways (sec(x))'=sec(x)tan(x)

myininaya (myininaya):

now you just replace the x's in sec(x)tan(x) with pi/3 and evaluate

myininaya (myininaya):

use the unit circle

OpenStudy (anonymous):

thank you !